OpenStudy (anonymous):

If f(x)+x^5[f(x)]^3=2052 and f(2)=4, find f′(2)=

OpenStudy (anonymous):

f(x)+x^5[f(x)]^3 is that one or two equations??

OpenStudy (anonymous):

Oops i mean f(x)+x^5[f(x)]^3 =2052

OpenStudy (anonymous):

it's one equation

hartnn (hartnn):

have you tried differentiating the left side?? you'll need product rule, chain rule, \(\Large \dfrac{d}{dx} x^n = n x^{n-1}\) \(\Large \dfrac{d}{dx} f(x)^n = n f(x)^{n-1} f'(x)\)

OpenStudy (anonymous):

That's what I figured I was suppose to do, but I don't see how I would do that.